# A Benchmark for Long-Form Medical Question Answering

### v1

@@ 2 Data Preparation
Figure 2: The distribution of queries in conversations (excluding outliers).
@@ 2.3 Inter-Dataset Similarity Analysis
Figure 3: Similarity graph of consumer medical question answering datasets. Each node represents a dataset and the number on the edge shows the mean similarity score of a pair of datasets.
@@ 3.3 Human Evaluation
Table 2: Our pairwise comparison expert annotation scheme.
@@ 3.4 LLM-as-a-Judge
To design our LLM-as-a-judge prompt template we merged the pair-v2 system prompt in Zheng et al., [27], and the pairwise evaluation template in WildBench [12]. Our full prompt is shown in Figure 12. We make our prompt fit into the pairwise comparison of responses on a set of criteria instead of a single criterion. We ran two LLMs as our judges including gpt-4o-2024-05-13 and claude-3-5-sonnet-20240620.
@@ 3.4.1 Testing Robustness
Before using LLM judgments and comparing them with human evaluations, we ran an analysis to see how robust and consistent LLMs are in their judgment. We ran each LLM six times on each batch –three times with the default order of responses, Response A and B (we refer to the three votes from these three times as an ab run), and another three times with a reversed order of responses (votes referred to as a ba run) to ensure the LLM judgments were not affected by positional biases of responses. Out of all 4,800 runs [^11] There are 4 batches, each with 100 questions, and 6 criteria per question, ran by 2 LLM judges given by LLMs, there were 174 cases (%4) among ab runs and 199 cases (%4) in ba runs where the LLM was not consistent in its judgment. And these inconsistencies are all from the gpt-4o model. Before moving to the next step, we resolved the inconsistencies by finding the majority vote for each run. There were only 3 cases in total where there was not a majority vote. In those cases, since tie was among the votes, we chose that as the majority.
After finding the majority votes for each run, we compared the ab and ba run votes for each model individually. We found that in 806 cases (%17 of 4,800 total runs, 397 for gpt-4o and 409 for claude) there’s a disagreement in model judgments between ab and ba runs. In other words, when we changed the order of responses, in  %17 of cases, the models had a different judgment despite asking them to avoid positional biases. The highest disagreements were related to the efficiency, correctness, and reasoning criteria, and the least disagreement was for the bias criterion with only one case and only for the claude model (details can be found in Appendix A.2.) Following Zheng et al., [27], we took a conservative approach and set the vote as tie and neither (for harmfulness and bias criteria) to resolve the disagreements. Then we compared gpt-4o and claude judgments and measured the correlation among their votes. There’s %96 percentage and %94 Cohen’s kappa agreement between the two models with chance agreement being %33. In the end, we resolve the disagreement between the two models based on the conservative approach described earlier and consider one final vote as our LLM-as-a-judge vote.
Figure 4: The distribution of the cumulative number of human votes and LLMs judgments across four annotation batches.
@@ 4 Results and Analysis
Turning to LLM-as-a-judge results and their comparison with human evaluations, we find general agreement across all batches and criteria. However, there remains a noticeable gap in the alignment between LLM votes and human labels.
@@ 4.1 Annotator Agreement
Table 3: Annotator agreement on annotation batches each with 100 examples. In each batch, we report the percentage (observed) agreement shown as P and chance agreement shown as C for each criterion.
@@
Table 4: Annotator disagreement counts on annotation batches across evaluation criteria.
@@ A.2 LLM-as-a-Judge Statistics
Table 5: Number of disagreements between ab and ba runs for each model based on the evaluation criteria.
@@
Correctness 81 78
Helpfulness 72 73
Harmfulness 71 78
Reasoning 79 79
Efficiency 94 100
Bias - 1
@@ Appendix B Inference Endpoints
Figure 9: The prompt template for annotating whether a query asks a direct medical or health-related question.
Figure 10: The prompt template for generating answers for medical questions. There’s a minor difference between GPT-4/Claude’s prompt at the end to accommodate output response in JSON format
Figure 11: The prompt template for annotating difficulty level of medical questions.

### v2

@@ 2 Data Preparation
Figure 2: The distribution of queries in conversations (excluding outliers)
@@ 2.3 Inter-Dataset Similarity Analysis
Figure 3: Similarity graph of consumer medical question answering datasets
@@ 3.3 Human Evaluation
Table 2: Our pairwise comparison expert annotation scheme
@@ 3.4 LLM-as-a-Judge
To design our LLM-as-a-judge prompt template we merged the pair-v2 system prompt in Zheng et al., [27], and the pairwise evaluation template in WildBench [12]. Our full prompt is shown in Figure 12. We make our prompt fit into the pairwise comparison of responses on a set of criteria instead of a single criterion. We ran two LLMs as our judges including gpt-4o-2024-08-06 and claude-3-5-sonnet-20241022.
@@ 3.4.1 Testing Robustness
Before using LLM judgments and comparing them with human evaluations, we conducted an analysis to assess the robustness and consistency of LLM judgments. Each LLM was run six times on each batch: three times with the default order of responses, Response A and B, and three times with the reversed order of responses. This was done to ensure that LLM judgments were not affected by positional biases of responses. We refer to the votes from the first three runs with the default order of responses as ab runs, and the votes from the three runs with the reversed order of responses as ba runs. Out of all 4,800 runs1111114 batches $\times$× 100 questions $\times$× 6 criteria per question $\times$× 2 LLM judges = 4,800 given by LLMs, there were 191 cases (%4) among ab runs and 213 cases (%4) in ba runs where the LLM was not consistent in its judgment. These inconsistencies in the gpt-4o model are approximately two times those in the claude model. Before moving to the next step, we resolved the inconsistencies by finding the majority vote for each run type.
After finding the majority votes, we compared the ab and ba run votes for each model individually. We found that in 811 cases (%17 of 4,800 total runs, 394 for gpt-4o and 417 for claude) there’s a disagreement in model judgments between ab and ba runs. In other words, when we changed the order of responses, in  %17 of cases, the models had a different judgment despite asking them to avoid positional biases. The highest disagreements were related to the efficiency and correctness criteria, and the least disagreement was for the bias criterion (details can be found in Appendix A.2.) Following Zheng et al., [27], we took a conservative approach and set the vote as tie and neither (for harmfulness and bias criteria) to resolve the disagreements. Then we compared gpt-4o and claude judgments and measured the correlation among their votes. There’s %71 percentage and %55 Cohen’s kappa agreement between the two models with chance agreement being %37. In the end, we resolve the disagreement between the two models based on the conservative approach described earlier and consider one final vote as our LLM-as-a-judge vote.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Medical DoctorsLLM-as-a-Judge
Figure 4: Distribution of votes by medical doctors (left) and LLM judge (right)
@@ 4 Results and Analysis
Turning to LLM-as-a-judge results and their comparison with human evaluations, we find general agreement across all batches and criteria. However, there remains a gap in the alignment between LLM votes and human labels.
@@ 4.1 Annotator Agreement
Table 3: Annotator agreement on annotation batches. P: Percentage (observed) agreement; C: Chance agreement
@@
Table 4: Annotator disagreement counts on annotation batches across evaluation criteria
@@ A.2 LLM-as-a-Judge Statistics
Table 5: Count of disagreements between ab and ba runs by model and evaluation criteria
@@
Correctness 72 76
Helpfulness 77 63
Harmfulness 54 38
Reasoning 88 52
Efficiency 103 181
Bias - 7
@@ Appendix B Inference Endpoints
Figure 9: The prompt template for annotating whether a query asks a direct medical or health-related question
Figure 10: The prompt template for generating answers for medical questions. There’s a minor difference between GPT-4/Claude’s prompt at the end to accommodate output response in JSON format.
Figure 11: The prompt template for annotating difficulty level of medical questions